# algorithmic

### v2

@@ 6.3 Multi-event prediction
We now consider the multi-event case ($K>1$), where the $K$ events are not mutually-exclusive and more than one event can happen per instance. We evaluate MENSA and state-of-the-art baselines in the MIMIC-IV, Rotterdam, PRO-ACT and EBMT test sets. Table 5 shows the prediction results. Overall, MENSA has the best (highest) global local concordance index in the Rotterdam and PRO-ACT datasets, and the best (highest) local concordance index in the MIMIC-IV, Rotterdam and PRO-ACT datasets. This shows a clear advantage in the proposed joint optimization strategy in MENSA, whereas the rival models, except the Hierarchical model, learn a separate model for each event and thus cannot leverage any shared information. MENSA also generally give accurate predictions of the survival function in terms of the integrated Brier Score, and accurate time-to-event estimates in terms of the margin MAE.
@@
Table 5: Multi-event prediction results. We report the mean ($\pm$ SD.) averaged over 10 experiments. D-calibration counts the number of experiments in which the respective model was D-calibrated. To improve readability, we have multiplied the CI and IBS results by 100, and divided the mMAE results in MIMIC-IV by 100.
As a practical motivation for the multi-event case, we consider the problem of predicting functional decline in patients suffering from ALS – which is a rapidly progressing motor neuron disease with a medium survival time of 3 to 5 years. However, instead of simply predicting the time to death (Kuan et al., 2023), we use our method to predict the time to functional decline in patients, which in our case consists of four distinct but related events: “Speaking”, “Swallowing”, “Handwriting” and “Walking” (see Section 5.2 and Appendix D for details). Such predictions can enable medical doctors to make informed, personalized decisions by tailoring treatments and interventions to an individual’s predicted trajectory. This is particularly critical in ALS, where timely interventions – e.g., assistive devices, nutritional support, and respiratory care – can greatly improve a patient’s quality of life. With MENSA, we can train a single survival model to predict an individual’s disease trajectory as $K$ individual survival distributions (ISDs). Figure 11 shows such ISDs, from baseline ($t=0$) to 500 days later, for the $i$-th patient in the PRO-ACT dataset using different numbers of mixture components ($\Psi$) for each event. When $\Psi=1$, the curves exhibit proportional hazards between them. When $\Psi>1$, the curves have more degrees of freedom and can thus model non-proportional hazards in this scenario. However, this also increases the number of parameters for each event, which comes with some pitfalls – e.g., the risk of overfitting and additional computational costs.
@@ 6.4 Model complexity
Deep learning survival models are usually applied for high-dimensional tasks, where the number of input parameters is large, so parameter and memory efficiency becomes important in such cases. We thus compare our model to rival baselines in terms of the number of trainable parameters and the number of floating-point operations per second (FLOPs) as a function of the number of features. Figure 12 presents the empirical results in single-event, competing-risks and multi-event scenarios. The number of FLOPs is calculated using the “fvcore” library [^2] Meta Research’s fvcore library, available at https://github.com/facebookresearch/fvcore (December 8, 2024).. In the single-event case, MENSA has a slower increase in parameters and FLOPs as the number of covariates increases. It is similar to other models like DSM and Deepsurv in terms of the rate of increase, but requires fewer parameters than DeepHit and MTLR. In competing risks, both the DSM and MENSA models exhibit relatively slow growth in terms of parameters and FLOPs, making them computationally-efficient for large datasets (more than 100 features). Lastly, in the multi-event case, models that cannot learn events jointly have to be trained a total of $K$ times to make predictions for $K$ events (see Table 6). In this case, the number of parameters and FLOPs in MENSA grow more steadily, compared to other models, like DeepSurv or DeepHit, and the growth rate does not accelerate drastically, like in the Hierarchical model. Based on these observations, MENSA is more computationally-efficient in terms of the number of parameters and the number of TFLOPs relative to other models, e.g., Hierarchical, DeepSurv and MTLR. This is further exemplified in multi-event scenarios, which in many instances require a per-event model to be trained.
@@ 7.1 Shared feature layer
We perform an ablation study to validate the effectiveness of the proposed shared feature layer in MENSA, $\Phi(x)$. The shared layer takes as input the features $\bm{x}^{(i)}$ and outputs a vector, $\tilde{\bm{x}}^{(i)}$, that captures the latent representation that is common to the events. Presumably, this allows the model to leverage shared information among the events, and shared feature layers have previously been used in competing-risks and multi-event works (Lee et al., 2018; Kim et al., 2021; Tjandra et al., 2021). We first train MENSA with a shared feature layer across all events, followed by training it with a separate feature layer specific to each event. Table 7 presents the results. In the PRO-ACT dataset, a shared layer gives small improvements in the global concordance index (70.74 vs. 70.68), but a small degradation in the local concordance index (76.12 vs. 76.36). This is possibly due to the lack of conditional dependencies in the PRO-ACT dataset. In the Rotterdam dataset, we see a minor improvement in margin MAE (18.91 vs. 18.71) using a shared layer and a notable improvement in the local concordance index (88.70 vs. 87.96).
@@ 8 Discussion
Limitations. One limitation to the present work is that the proposed method assumes conditional independent censoring, i.e., the event time is independent of the censoring time, given the patient’s features ($\bm{x}^{(i)}$). In other words, it assumes that any relationship between event and censoring times is captured in the features. However, this assumption may fail to fully account for the dependency if important covariates are missing in the analysis. To this end, we evaluated MENSA under dependent censoring in the single-event case and measured its prediction bias using the Survival-$\ell_{1}$, i.e., the difference between the predicted survival curve and the true survival curve. We found that the bias was steady in the linear case under dependence from a Frank or Clayton copula, and well-below established baselines, but increased as a function of Kendall’s $\tau$ in the nonlinear case, more aligned with the baselines. This shows that MENSA does make biased survival predictions in the nonlinear case in the presence of dependent censoring, although the incurred prediction bias was lower in MENSA than in several rival models.
Another limitation, based on our observations, is that parametric survival models, such as MENSA, often require more data to train effectively and achieve good results compared to semiparametric models, such as Cox-based regression models. These models make fewer assumptions about the underlying survival time distribution, as they only model the hazard function’s proportionality using e.g., the exponential distribution, and leave the baseline hazard unspecified. This makes the CoxPH model and Cox-based models, e.g., CoxBoost and DeepSurv, less sensitive to the dataset’s size and distribution assumptions. On the other hand, parametric models, e.g., DSM and MENSA, assume a specific distribution (e.g., Weibull, exponential) for the survival times. The flexibility of the model can lead to better performance if the parametric assumption is correct and data is plentiful, but it also adds more parameters that need to be estimated.
@@
Future work. The next natural step in our work is to address the aforementioned prediction bias under dependent censoring, as we demonstrated on the synthetic dataset in the single-event case. All survival models in this work, both the established baselines and MENSA, assume conditional independent censoring, and thus may give biased parameter estimates when dependent censoring is present (Emura and Chen, 2018, Ch. 3). Although we can reduce dependent censoring simply by collecting more training data (i.e., features) (Emura and Chen, 2018, Ch. 1), ethical or practical constraints in medical studies may prevent us from doing so, especially in large-scale or long-term studies. Alternatively, we can incorporate the dependence between event distributions outside the features in the parameter estimation by learning it as a function of the data: Foomani et al. (2023) proposed learning a survival model under dependent censoring by modifying the survival likelihood in Equation 1 to include a copula from the Archimedean family (e.g., Clayton, Frank). Although a detailed discussion is beyond the scope of this work, this copula-dependent likelihood is split into four terms: the first two terms account for the log-likelihood of observing an event and its conditional probability given the copula, while the last two terms do the same for the censoring. The copula’s partial derivatives have closed-form solutions for Archimedean copulas, which enable efficient optimization using gradient-based methods in MLE. This approach can be extended to a multi-event scenario by first estimating a $K$-variate copula model for the joint event distribution and then incorporate the model into the survival likelihood across $K$ events.
@@
An alternative approach is to parameterize the copula as a neural network, which eliminates the need to specify the copula family – e.g., Clayton or Frank – beforehand. Recent studies have adopted this approach and shown promising results for single-event (Zhang et al., 2024) and competing-risks (Liu et al., 2024) scenarios. Zhang et al. (2024) proposed DCSurvival, an end-to-end framework that estimates the survival marginals using monotonic neural density estimators (NDE) and parameterizes the (Archimedean) copula generator function as a neural network, $\varphi_{nn}$. Unlike Foomani et al. (2023), DCSurvival does not make assumptions about the copula family, other than that it is Archimedean, since it learns the copula nonparametrically. However, using a single Archimedean copula can characterize only a symmetric dependency relationship. To this end, Liu et al. (2024) proposed HACSurv, which extends DCSurvival to competing risks by modeling $C(u_{0},...,u_{K})$ using a flexible hierarchical Archimedean copula (HAC). Thus, this model can capture diverse forms of dependencies between competing events. An open avenue for future research is to extend these hierarchical copulas to the multi-event setting.
Boarder impacts. This paper proposes a new method for survival analysis, where there a multiple events of interest, accompanied by experiments on five publicly-available clinical datasets to show the practical usefulness of the model. These datasets were used following the guidance of the respective data providers and domain experts. We acknowledge that survival analysis can give valuable insights in treatment planning or personalized care, but can also have adverse impacts if misinterpreted or misused, or if the dataset contains systemic biases, leading to inequities or discrimination. The deployment of such models comes with several potential risks: (1) If the training data is not representative of the population and subgroups the models are used within, they may disproportionately favor certain groups over others. (2) Covariate-based survival models, similar to other machine learning models, can unintentionally use sensitive variables (e.g., race, gender, or socioeconomic status), which can lead to biased predictions. (3) Survival models can replace human decision-makers in certain domains, such as healthcare, and autonomously decide who gets access to treatment and who does not based on some criteria. This, again, may reproduce or amplify existing biases in the data, leading to unfair or discriminatory results. To mitigate these risks, we emphasize the importance of adequate evaluation and testing before deploying such models, and that these models should be used as a decision-support tool in the healthcare section rather than as an independent oracle for decision-making.
@@ 9 Conclusion
We have presented MENSA, a novel method for multi-event survival analysis. Given an instance, the model predicts the time until that instance experiences each of several different events, including being event-free. In real clinical datasets, we show that the proposed method led to improvements over well-accepted baselines in several performance metrics, especially in terms of predictive accuracy (Brier Score and MAE), local discrimination (i.e., ranking of events within individuals) and distribution calibration. An ablation study showed that combining the traditional log-likelihood in multi-event with a trajectory-based likelihood, which promotes learning temporal patterns between events, consistently improved the Brier Score and MAE in two multi-event datasets. We also highlighted the advantages that our model holds over state-of-the-art baselines in terms of computational efficiency and resource utilization, especially in the multi-event scenario, where traditional survival models need to train a model for each event. As a practical motivation for the multi-event case, we predicted the time to functional decline in ALS patients, which in our case consisted of four separate, but related events. This application can facilitate the design of personalized treatment plans for the disease and improve a patient’s quality of life.
Acknowledgments
This research received support from the Natural Science and Engineering Research Council of Canada (NSERC), the Canadian Institute for Advanced Research (CIFAR), and the Alberta Machine Intelligence Institute (Amii).
@@ Appendix B Evaluation metrics
Harrell’s CI: The concordance index (CI) measures the discriminative performance of a survival model by calculating the proportion of concordant pairs among all comparable pairs. A pair is considered comparable if we can determine who has the event first. It is defined as (Harrell Jr. et al., 1982):
@@
Global CI: The global CI is calculated as the average of the CI values across all specific events, providing an overall measure of a model’s discriminative performance for all competing and multiple events. This approach follows previous work (Lee et al., 2018; Katzman et al., 2018), ensuring a comprehensive evaluation of the model’s global discrimination ability. It is defined as:
@@
Local CI: The local CI is computed by averaging CI scores across multiple events for each instance, and was proposed by Tjandra et al. (2021). This metric evaluates the model’s ability to discriminate between multiple events within a single instance. It is defined as (Tjandra et al., 2021):
@@
BS/IBS: The Brier Score (BS) is defined as the mean squared difference between the predicted survival curve and the Heaviside step function of the observed event. The integrated Brier Score (Graf et al., 1999) (IBS) aggregates the Brier Scores across multiple time points to provide a single measure of model performance. We use inverse probability weighting (IPCW) to handle censored events. The BS is defined as:
@@
MAE: The mean absolute error (MAE) is the absolute difference between the predicted and actual survival times. Given an individual survival distribution, $S\left(t\mid\bm{x}_{i}\right)=\text{Pr}\left(T>t\mid\bm{x}_{i}\right)$, we calculate the predicted survival time $\hat{t}^{(i)}$ using the median time of the predicted survival distribution (Qi et al., 2023b):
@@
For censored individuals, we calculate the marginal MAE (mMAE), as proposed by Haider et al. (2020):
@@
D-calibration: Distribution Calibration (Haider et al., 2020) measures the calibration performance of $S(t)$, expressing to what extent the predicted probabilities can be trusted. We assess this using a Pearson’s ${\chi}^{2}$ goodness-of-fit test. For any probability interval $[a,b]\in[0,1]$, we define $D_{m}(a,b)$ as the group of individuals in the dataset $D$ whose predicted probability of event is in the interval $[a,b]$ (Qi et al., 2023a). A model is D-calibrated if the proportion of individuals $|D_{m}(a,b)|/|D|$ is statistically similar to the amount $b-a$. In the results, we count the number of times a model is D-calibrated (at $p>0.05$) for each event in 10 experiments.
Survival-$\ell_{1}$: The Survival-$\ell_{1}$ is the L1 distance between the ground-truth survival curve, $S_{T\mid X}$, and the estimated survival curve, $\hat{S}_{T\mid X}$, over the event horizon (Foomani et al., 2023). We normalize the area between the survival curves by $T_{\text{max}}^{(i)}=S_{T\mid X^{(i)}}^{-1}\left(Q\|\cdot\|\right)$ to ensure that the duration spanned by an instance’s survival curve does not influence that instance’s contribution to the metric relative to other instances. The Survival-$\ell_{1}$ is defined as (Foomani et al., 2023):

### v3

@@ 6.3 Multi-event prediction
We now consider the multi-event case ($K>1$), where the $K$ events are not mutually exclusive and more than one event can happen per instance. We evaluate MENSA and state-of-the-art baselines in the MIMIC-IV, Rotterdam, PRO-ACT and EBMT test sets. Table 5 shows the prediction results. Overall, MENSA has the best (highest) global local concordance index in the Rotterdam and PRO-ACT datasets, and the best (highest) local concordance index in the MIMIC-IV, Rotterdam and PRO-ACT datasets. This shows a clear advantage in the proposed joint optimization strategy in MENSA, whereas the rival models, except the Hierarchical model, learn a separate model for each event and thus cannot leverage any shared information. MENSA also generally give accurate predictions of the survival function in terms of the integrated Brier Score, and accurate time-to-event estimates in terms of the margin MAE.
@@
Table 5: Multi-event prediction results. We report the mean ($\pm$ SD.) averaged over 10 experiments. D-calibration counts the number of experiments in which the respective model was D-calibrated. To improve readability, CI and IBS were multiplied by 100, and mMAE in MIMIC-IV was divided by 100.
As a practical motivation for the multi-event case, we consider the problem of predicting functional decline in patients suffering from ALS – which is a rapidly progressing motor neuron disease with a medium survival time of 3 to 5 years. However, instead of simply predicting the time to death [28], we use our method to predict the time to functional decline in patients, which in our case consists of four distinct but related events: “Speaking”, “Swallowing”, “Handwriting” and “Walking” (see Section 5.2 and Appendix D for details). Such predictions can enable medical doctors to make informed, personalized decisions by tailoring treatments and interventions to an individual’s predicted trajectory. This is particularly critical in ALS, where timely interventions – e.g., assistive devices, nutritional support, and respiratory care – can greatly improve a patient’s quality of life. With MENSA, we can train a single survival model to predict an individual’s disease trajectory as $K$ individual survival distributions (ISDs). Figure 13 shows such ISDs, from baseline ($t=0$) to 500 days later, for the $i$-th patient in the PRO-ACT dataset using different numbers of mixture components ($\Psi$) for each event. When $\Psi=1$, the curves exhibit proportional hazards between them. When $\Psi>1$, the curves have more degrees of freedom and can thus model non-proportional hazards in this scenario. However, this also increases the number of parameters for each event, which comes with some pitfalls – e.g., the risk of overfitting and additional computational costs.
@@ 6.4 Model complexity
Deep learning survival models are usually applied for high-dimensional tasks, where the number of input parameters is large, so parameter and memory efficiency becomes important in such cases. We thus compare our model to rival baselines in terms of the number of trainable parameters and the number of floating-point operations per second (FLOPs) as a function of the number of features. Figure 14 shows the results wrt model complexity in the single-event, competing-risks and multi-event scenarios. The number of FLOPs is calculated using the “fvcore” library [^2] Meta Research’s fvcore library, available at https://github.com/facebookresearch/fvcore (December 8, 2024).. In the single-event case, MENSA has a slower increase in parameters and FLOPs as the number of covariates increases. It is similar to other models like DSM and Deepsurv in terms of the rate of increase, but requires fewer parameters than DeepHit and MTLR. In competing risks, both the DSM and MENSA models exhibit relatively slow growth in terms of parameters and FLOPs, making them computationally-efficient for large datasets (more than 100 features). Lastly, in the multi-event case, models that cannot learn events jointly have to be trained a total of $K$ times to make predictions for $K$ events (see Table 6). In this case, the number of parameters and FLOPs in MENSA grow more steadily, compared to other models, like DeepSurv or DeepHit, and the growth rate does not accelerate drastically, like in the Hierarchical model. Based on these observations, MENSA is more computationally-efficient in terms of the number of parameters and the number of FLOPs relative to other models, e.g., Hierarchical, DeepSurv and MTLR. This is further exemplified in multi-event scenarios, which in many instances require a per-event model to be trained.
@@ 7.1 Shared feature layer
We perform an ablation study to validate the effectiveness of the proposed shared feature layer in MENSA, $\Phi(x)$. The shared layer takes as input the features $\bm{x}^{(i)}$ and outputs a vector, $\tilde{\bm{x}}^{(i)}$, that captures the latent representation that is common to the events. Presumably, this allows the model to leverage shared information among the events, and shared feature layers have previously been used in competing-risks and multi-event works [30, 24, 51]. We first train MENSA with a shared feature layer across all events, followed by training it with a separate feature layer specific to each event. Table 7 presents the results. In the PRO-ACT dataset, a shared layer gives small improvements in the global concordance index (70.74 vs. 70.68), but a small degradation in the local concordance index (76.12 vs. 76.36). This is possibly due to the lack of conditional dependencies in the PRO-ACT dataset. In the Rotterdam dataset, we see a minor improvement in margin MAE (18.91 vs. 18.71) using a shared layer and a notable improvement in the local concordance index (88.70 vs. 87.96).
@@ 8 Discussion
Limitations. This work has limitations. First, the proposed method assumes conditional independent censoring, i.e., the event time is independent of the censoring time, given the patient’s features. In other words, it assumes that any relationship between event and censoring times is captured in the features. However, this assumption may fail to fully account for the dependency if important covariates are missing in the analysis. To this end, we evaluated MENSA under dependent censoring in the single-event case and measured its prediction bias using the Survival-$\ell_{1}$, i.e., the difference between the predicted survival curve and the true survival curve. We found that the bias was steady in the linear case under dependence from a Frank or Clayton copula, and well-below established baselines, but increased as a function of Kendall’s $\tau$ in the nonlinear case, more aligned with the baselines. This shows that MENSA does make biased survival predictions in the nonlinear case in the presence of dependent censoring, although the incurred prediction bias was lower in MENSA than in several rival models. Second, parametric survival models, based on our observations, often require more data to train effectively and achieve good results compared to semiparametric models, such as Cox-based regression models. These models make fewer assumptions about the underlying survival time distribution, as they only model the hazard function’s proportionality using e.g., the exponential distribution, and leave the baseline hazard unspecified. This makes the CoxPH model and Cox-based models, e.g., CoxBoost and DeepSurv, less sensitive to the dataset’s size and distribution assumptions. On the other hand, parametric models, e.g., DSM and MENSA, assume a specific distribution (e.g., Weibull, exponential) for the survival times. The flexibility of the model can lead to better performance if the parametric assumption is correct and data is plentiful, but it also adds more parameters that need to be estimated.
@@
Future work. The next natural step in our work is to address the aforementioned prediction bias under dependent censoring, as we demonstrated on the synthetic dataset in the single-event case. All survival models in this work, both the established baselines and MENSA, assume conditional independent censoring, and thus may give biased parameter estimates when dependent censoring is present [8, Ch. 3]. Although we can reduce dependent censoring simply by collecting more training data (i.e., features) [8, Ch. 1], ethical or practical constraints in medical studies may prevent us from doing so, especially in large-scale or long-term studies. Alternatively, we can incorporate the dependence between event distributions outside the features in the parameter estimation by learning it as a function of the data: Foomani et al. [12] proposed learning a survival model under dependent censoring by modifying the survival likelihood in Equation 15 to include a copula from the Archimedean family (e.g., Clayton, Frank). Although a detailed discussion is beyond the scope of this work, this copula-dependent likelihood is split into four terms: the first two terms account for the log-likelihood of observing an event and its conditional probability given the copula, while the last two terms do the same for the censoring. The copula’s partial derivatives have closed-form solutions for Archimedean copulas, which enable efficient optimization using gradient-based methods in MLE. This approach can be extended to a multi-event scenario by first estimating a $K$-variate copula model for the joint event distribution and then incorporate the model into the survival likelihood across $K$ events.
@@
An alternative approach is to parameterize the copula as a neural network, which eliminates the need to specify the copula family – e.g., Clayton or Frank – beforehand. Recent studies have adopted this approach and shown promising results for single-event [56] and competing-risks [34] scenarios. Zhang et al. [56] proposed DCSurvival, an end-to-end framework that estimates the survival marginals using monotonic neural density estimators (NDE) and parameterizes the (Archimedean) copula generator function as a neural network, $\varphi_{nn}$. Unlike Foomani et al. [12], DCSurvival does not make assumptions about the copula family, other than that it is Archimedean, since it learns the copula nonparametrically. However, using a single Archimedean copula can characterize only a symmetric dependency relationship. To this end, Liu et al. [34] proposed HACSurv, which extends DCSurvival to competing risks by modeling $C(u_{0},...,u_{K})$ using a flexible hierarchical Archimedean copula (HAC). Thus, this model can capture diverse forms of dependencies between competing events. An open avenue for future research is to extend these hierarchical copulas to the multi-event setting.
@@ 9 Conclusion
We have presented MENSA, a novel method for multi-event survival analysis. Given an instance, the model predicts the time until that instance experiences each of several different events, including being event-free. In real clinical datasets, we showed that MENSA led to improvements over well-accepted baselines in several performance metrics, especially for Harrell’s concordance index (within events and within individuals), Brier Score and MAE. An ablation study showed that combining the traditional log-likelihood in multi-event with a trajectory-based likelihood, which promotes learning temporal patterns between events, consistently improved the Brier Score and MAE in two multi-event datasets. We also highlighted the advantages that our model holds over state-of-the-art baselines in terms of computational efficiency and resource utilization, especially in the multi-event scenario, where traditional survival models need to train a model for each event. As a practical motivation for the multi-event case, we predicted the time to functional decline in ALS patients, which in our case consisted of four separate, but related events. This application can facilitate the design of personalized treatment plans for the disease and improve a patient’s quality of life.
@@
Acknowledgements
[…]
Declarations
Conflict of interest
The authors declare that they have no conflict of interest.
Ethics approval
Ethics approval is not applicable. The study did no human trials and all data were completely anonymized.
Funding
R.G received support from the Natural Science and Engineering Research Council of Canada (NSERC), the Canadian Institute for Advanced Research (CIFAR), and the Alberta Machine Intelligence Institute (Amii).
Data availability
All datasets used in this study are publicly available.
Code availability
The code is publicly available at https://github.com/thecml/mensa.
Consent for publication
Not applicable.
Materials availability
Not applicable.
Author contribution
Not applicable.
@@ Appendix B Evaluation metrics
Harrell’s CI: The concordance index (CI) measures the discriminative performance of a survival model by calculating the proportion of concordant pairs among all comparable pairs. A pair is considered comparable if we can determine who has the event first. It is defined as [17]:
@@
Global CI: The global CI is calculated as the average of the CI values across all specific events, providing an overall measure of a model’s discriminative performance for all competing and multiple events. This approach follows previous work [30, 23], ensuring a comprehensive evaluation of the model’s global discrimination ability. It is defined as:
@@
Local CI: The local CI is computed by averaging CI scores across multiple events for each instance, and was proposed by Tjandra et al. [51]. This metric evaluates the model’s ability to discriminate between multiple events within a single instance. It is defined as [51]:
@@
BS/IBS: The Brier Score (BS) is defined as the mean squared difference between the predicted survival curve and the Heaviside step function of the observed event. The integrated Brier Score [14] (IBS) aggregates the Brier Scores across multiple time points to provide a single measure of model performance. We use inverse probability weighting (IPCW) to handle censored events. The BS is defined as:
@@
MAE: The mean absolute error (MAE) is the absolute difference between the predicted and actual survival times. Given an individual survival distribution, $S\left(t\mid\bm{x}_{i}\right)=\text{Pr}\left(T>t\mid\bm{x}_{i}\right)$, we calculate the predicted survival time $\hat{t}^{(i)}$ using the median time of the predicted survival distribution [41]:
@@
For censored individuals, we calculate the marginal MAE (mMAE), as proposed by Haider et al. [16]:
@@
D-calibration: Distribution Calibration [16] measures the calibration performance of $S(t)$, expressing to what extent the predicted probabilities can be trusted. We assess this using a Pearson’s ${\chi}^{2}$ goodness-of-fit test. For any probability interval $[a,b]\in[0,1]$, we define $D_{m}(a,b)$ as the group of individuals in the dataset $D$ whose predicted probability of event is in the interval $[a,b]$ [40]. A model is D-calibrated if the proportion of individuals $|D_{m}(a,b)|/|D|$ is statistically similar to the amount $b-a$. In the results, we count the number of times a model is D-calibrated (at $p>0.05$) for each event in 10 experiments.
Survival-$\ell_{1}$: The Survival-$\ell_{1}$ is the L1 distance between the ground-truth survival curve, $S_{T\mid X}$, and the estimated survival curve, $\hat{S}_{T\mid X}$, over the event horizon [12]. We normalize the area between the survival curves by $T_{\text{max}}^{(i)}=S_{T\mid X^{(i)}}^{-1}\left(Q\|\cdot\|\right)$ to ensure that the duration spanned by an instance’s survival curve does not influence that instance’s contribution to the metric relative to other instances. The Survival-$\ell_{1}$ is defined as [12]: